# Integral bases, perfect matchings, and the Petersen graph

### v2

@@ 5 Proof of the Intersection Theorem
We first reduce to the case where $G$ is a brick. To this end, suppose $G$ has a tight cut, and let $G_{1},G_{2}$ be the corresponding cut-contractions, where $G_{1}$ is bipartite, and $G_{2}$ is a Petersen-free near-brick. We know that $G_{2}$ is non-BvN, by 4.3. By the induction hypothesis, $G_{2}$ has a perfect matching $N$ and a separating facet-defining cut $C$ such that $|N\cap C|=3$. Then $C$ is a separating facet-defining cut of $G$, by 4.7. Thus, by extending $N$ to a perfect matching $M$ of $G$, we see that $(M,C)$ is a desired pair.
@@
Otherwise, $H\neq G\setminus e$, so $G\setminus e$ has a tight cut. Every tight cut of $G\setminus e$ is a barrier cut of the form $\delta_{G\setminus e}(X)$ with barrier $B\subset\bar{X}$, by 4.6. Observe that $P_{e}(G)\subseteq P(G;\delta_{G}(X))$, and since $P(G;\delta_{G}(X))\neq P(G)$ as $G$ is a brick, we must have $P_{e}(G)=P(G;\delta_{G}(X))$. A simple counting argument implies that $e$ joins either two singleton components of $G\setminus e\setminus B$, or it joins a singleton component to $X$. This implies that there exist either one or two maximal barriers only, summarized as follows.
@@
It can be shown through an elementary though terse argument that in each of Cases 1-3, there exist a $5$-cycle of $H$ with vertex set $Y$, and perfect matchings $M,M^{\prime}$ of $G$ containing $e$, such that $Y$ does not use contraction vertices of $H$, $Y$ is not incident to either ends of $e$, $|M\cap\delta_{G}(Y)|=3$, and $|M^{\prime}\cap\delta_{G}(Y)|=1$. [^2] For a detailed analysis, we refer the reader to the proof of ([5], Theorem 5.4), or ([12], Theorem 13.6). Moving on, we claim that $\delta_{G}(Y)$ is a separating facet-defining cut of $G$. Clearly, $\delta_{G\setminus e}(Y)$ is separating in $G\setminus e$ as it is so in $H$. Thus, as $e\in M^{\prime}$ and $|M^{\prime}\cap\delta_{G}(Y)|=1$, it follows that $\delta_{G}(Y)$ is separating in $G$. To see that $\delta_{G}(Y)$ is facet-defining in $G$, pick perfect matchings $M_{1},\ldots,M_{d}$ whose incidence vectors lie inside the facet $P_{e}(G)$ and are linearly independent; we may pick the first d1d-1 perfect matchings to have intersection $1$ with $\delta_{G}(Y)$, and the last one with intersection $5$, by 4.9. We then swap $M_{d}$ with the perfect matching $M^{\prime}\ni e$ which satisfies $|M^{\prime}\cap\delta(Y)|=1$, in order to obtain $d$ linearly independent vectors $M_{1},\ldots,M_{d-1},M^{\prime}$ in the face $P(G;\delta_{G}(Y))$, implying in turn that $\delta_{G}(Y)$ defines a facet of $P(G)$. Consequently, $(M,\delta_{G}(Y))$ is a desired pair.
@@
In the remaining case, $P(G;C)$ is a (d2)(d-2)-face of $P(G)$, contained in the facet $P_{e}(G)$. Consider now the other facet of $P(G)$ containing $P(G;C)$. It cannot be exposed by a non-negativity inequality, given that $P(G;C)$ is not contained in $\{x:x_{f}=0\}$ for any $f\in E\setminus e$. In particular, the other facet is exposed by a separating facet-defining cut of $G$, say $D$. In particular, $D\setminus e$ defines the same facet as $C\setminus e$, and so $D\setminus e,C\setminus e$ must be equivalent cuts as $G\setminus e$ is a near-brick, by 4.8. Thus, $|M\cap D|=|M\cap C|=3$, implying in turn that $(M,D)$ is a desired pair.
@@
Suppose some (d2)(d-2)-face of $P(G)$ is not exposed by a non-negativity inequality. Take a facet containing this face, which is exposed by a separating facet-defining cut $C=\delta(X)$. Then both $C$-contractions of $G$ are near-bricks, by 4.5, part (3). Given that the polytope $P(G;C)$ has a facet not exposed by a non-negativity inequality, one of the $C$-contractions, say $G/\bar{X}$, must be non-BvN, by 4.3.
We may assume that $G/\bar{X}$ is a non-BvN brick. To argue this, note first that the unique brick of $G/\bar{X}$ is non-BvN, by 4.3. If $G/\bar{X}$ is not this brick, then choose a minimal vertex subset $Y\subset X$ such that $\delta(Y)$ is a tight cut of $G/\bar{X}$. In particular, $P(G;C)\subseteq P(G;\delta(Y))$. As $G$ is a brick, $\delta(Y)$ is not tight in $G$, so $P(G;C)=P(G;\delta(Y))$, therefore $C,\delta(Y)$ are equivalent cuts and $G/Y/\bar{X}$ is bipartite, by 4.8. This, together with our minimal choice of $Y$, implies that $G/\bar{X}/\bar{Y}=G/\bar{Y}$ is a brick. Subsequently, by changing $C=\delta(X)$ and without changing the corresponding facet, if necessary, we may assume $G/\bar{X}$ is a non-BvN brick.
@@
    Suppose $Z\subset X$. Then $D$ is an odd cut of $G/\bar{X}$, so $R,D$ define the same facet of $P(G/\bar{X})$, so $R,D$ are equivalent cuts of $G/\bar{X}$ as $G/\bar{X}$ is a brick, by 4.8. Thus, $|D\cap N|=|R\cap N|=3$, so $|D\cap M|=|D\cap N|=3$, therefore $(M,D)$ is a desired pair.
@@
    Suppose $X\subset Z$. Then $(R,C,D)$ would be a cut triple defining a face triple, thus contradicting 4.10.
@@
    In the remaining case, $X$ and $Z$ cross. Let $I:=\delta(X\cap Z)$ and $U:=\delta(X\cup Z)$. Given that $P(G;C)\cap P(G;D)=P(G;R)\not\subseteq\{x:x_{e}=0\}$ for any $e\in E$, it follows from 4.2 that $x(C)+x(D)=x(I)+x(U)$ and $P(G;R)=P(G;C)\cap P(G;D)=P(G;I)\cap P(G;U)$.
    If $\{P(G;I),P(G;U)\}=\{P(G;R),P(G;C)\}$ or $\{P(G;R),P(G;D)\}$, then $(I,D,U)$ or $(I,C,U)$ would be a cut triple defining a face triple, respectively, thus contradicting 4.10.
    If $\{P(G;I),P(G;U)\}=\{P(G;R),P(G)\}$, then we have $P(G;I)=P(G;R)$ and $P(G;U)=P(G)$. This holds since otherwise $(R,C,U)$ would be a cut triple defining a face triple, thus contradicting 4.10. The equality $P(G;U)=P(G)$ implies that $U$ is a trivial cut as $G$ is a brick, while $P(G;I)=P(G;R)$ implies that $R,I$ are equivalent cuts in the brick $G/\bar{X}$, so $x(I)=x(R)$ for all $x\in P(G)$. Consequently, x(D) = x(I)+x(U)-x(C)=x(R)+1-x(C), so whenever $x(C)=1$, then $x(R)=x(D)$, so $|D\cap M|=|R\cap M|=3$, so $(M,D)$ is a desired pair.
    Otherwise, $\{P(G;I),P(G;U)\}=\{P(G;C),P(G;D)\}$, then we have $P(G;I)=P(G;D)$ and $P(G;U)=P(G;C)$. This holds since otherwise $(R,C,U)$ would be cut triple defining a face triple, thus contradicting 4.10. Now by replacing $D$ with $I$ we fall back to Case 1.
@@ 6 Proof of the Main Theorem
Suppose there is a tight cut $C$. Let $G_{1},G_{2}$ be the $C$-contractions of $G$, both of which are Petersen-free matching-covered graphs. By the induction hypothesis, for each $i\in[2]$, $\operatorname{lin}(P(G_{i}))$ has an integral basis $B^{i}$ that consists solely of the incidence vectors of some perfect matchings of $G_{i}$. Let $B:=B^{1}\odot B^{2}$ which is a subset of $\mathbb{Z}^{E}$ consisting of incidence vectors of some perfect matchings of $G$ that intersect $C$ just once, and is an integral basis for $\operatorname{lin}(P(G;C))$, by 4.1, part (3). Given that $C$ is a tight cut, it follows that $P(G;C)=P$, thus completing the induction step.
@@
Otherwise, $G$ is a non-BvN non-Petersen brick. By the Intersection Theorem, there exist a separating facet-defining cut $C=\delta(X)$ and a perfect matching $M$ such that $|C\cap M|=3$. Let $G_{1}:=G/X$ and $G_{2}:=G/\bar{X}$, both of which are near-bricks, by 4.5, part (3).
Suppose in the first case that both $G_{1},G_{2}$ are Petersen-free. Then by the induction hypothesis, for each $i\in[2]$, $\operatorname{lin}(P(G_{i}))$ has an integral basis $B^{i}$ that consists solely of the incidence vectors of some perfect matchings of $G_{i}$. Let $B^{\prime}:=B^{1}\odot B^{2}$ which is a subset of $\mathbb{Z}^{E}$ consisting of incidence vectors of some perfect matchings of $G$ that intersect $C$ just once, and is an integral basis for $\operatorname{lin}(P(G;C))$, by 4.1, part (3). We claim that $B:=B^{\prime}\cup\{\mathbf{1}_{M}\}$ is an integral basis for $\operatorname{lin}(P)$. Linear independence is clear as $|M\cap C|>1$ while $b(C)=1$ for all $b\in B^{\prime}$. Given that $\dim(P)=\dim(P(G;C))+1$, it therefore follows that $B$ is a basis for $\operatorname{lin}(P)$. It remains to prove that $B$ is an integral basis. To this end, let $z\in\operatorname{lin}(P)\cap\mathbb{Z}^{E}$, and express this vector as a unique linear combination of $B$: $z=\sum_{b\in B}\alpha_{b}b$. We need to show that $\alpha_{b}\in\mathbb{Z}$ for all $b\in B$.
@@
We may assume that the near-brick $G_{2}$ is a Petersen brick. If not, then choose a minimal subset $Z\subset X$ such that $\delta(Z)$ is a tight cut of $G/\bar{X}$. In particular, $P(G;C)\subseteq P(G;\delta(Z))$. As $G$ is a brick, $\delta(Z)$ is not tight in $G$, so $\delta(Z)$ and $C$ define the same facet of $P$, so they are equivalent cuts and $G/Z/\bar{X}$ is bipartite, by 4.8. This, together with our minimal choice of $Z$, implies that $G/\bar{Z}$ is a brick. Subsequently, by changing $C=\delta(X)$ to an equivalent cut if necessary, we may assume $G_{2}$ is a Petersen brick.
@@
To argue that $\delta(Y)$ is facet-defining for $G$, note first that the perfect matching $M$ can be redefined inside $G[X]$, if necessary, such that $|M\cap\delta(Y)|=1$. Secondly, by 4.1, part (4) and 4.9, there exist perfect matchings $M_{1},\ldots,M_{d}$ whose incidence vectors belong to $P(G;C)$ and form a basis for $\operatorname{lin}(P(G;C))$, where $|M_{1}\cap\delta(Y)|=5$, $|M_{i}\cap\delta(Y)|=1$ for $i\in\{2,\ldots,d\}$, and every edge of $G$ belongs to one of $M_{2},\ldots,M_{d}$. By swapping $M_{1}$ with $M$, we obtain $d$ perfect matchings $M,M_{2},\ldots,M_{d}$ whose incidence vectors belong to $P(G;\delta(Y))$ and are linearly independent, because $|M\cap C|=3$ while $|M_{i}\cap C|=1$ for $i\in\{2,\ldots,d\}$. Thus, $P(G;\delta(Y))$ is a facet of $P$.
@@
where each $A_{i}$ is the edge set of some $5$-cycle of $H_{i}$. Let $B:=B_{1}\odot B_{2}$, which clearly consists of some perfect matchings of $G$. By 4.1, part (2), $B$ is a lattice basis for $L$. Furthermore, we claim that the above set equalities imply
@@
$(\subseteq)$ is clear. $(\supseteq)$: Pick $f\in\bar{L}$ such that $f(A_{i})\equiv 0\pmod{2},~\forall i\in[p]$. Then $f=x\odot y$ for $x\in\bar{L}_{1}$ and $y\in\bar{L}_{2}$, where clearly $x(A_{i})\equiv 0\pmod{2},~i=1,\ldots,q$ and $y(A_{i})\equiv 0\pmod{2},~i=q+1,\ldots,p$. Thus, $x\in L_{1},y\in L_{2}$, so $f$ belongs to $L$, by 4.1, part (2). This completes the induction step. ∎
Acknowledgements

### v3

@@ 4 Proof of the Intersection Theorem
We first reduce to the case where $G$ is a brick. To this end, suppose $G$ has a tight cut, and let $G_{1},G_{2}$ be the corresponding cut-contractions, where $G_{1}$ is bipartite, and $G_{2}$ is a Petersen-free near-brick. We know that $G_{2}$ is non-BvN, by 3.3. By the induction hypothesis, $G_{2}$ has a perfect matching $N$ and a separating facet-defining cut $C$ such that $|N\cap C|=3$. Then $C$ is a separating facet-defining cut of $G$, by 3.7. Thus, by extending $N$ to a perfect matching $M$ of $G$, we see that $(M,C)$ is a desired pair.
@@
Otherwise, $H\neq G\setminus e$, so $G\setminus e$ has a tight cut. Every tight cut of $G\setminus e$ is a barrier cut of the form $\delta_{G\setminus e}(X)$ with barrier $B\subset\bar{X}$, by 3.6. Observe that $P_{e}(G)\subseteq P(G;\delta_{G}(X))$, and since $P(G;\delta_{G}(X))\neq P(G)$ as $G$ is a brick, we must have $P_{e}(G)=P(G;\delta_{G}(X))$. A simple counting argument implies that $e$ joins either two singleton components of $G\setminus e\setminus B$, or it joins a singleton component to $X$. This implies that there exist either one or two maximal barriers only, summarized as follows.
@@
It can be shown through an elementary though terse argument that in each of Cases 1-3, there exist a $5$-cycle of $H$ with vertex set $Y$, and perfect matchings $M,M^{\prime}$ of $G$ containing $e$, such that $Y$ does not use contraction vertices of $H$, $Y$ is not incident to either ends of $e$, $|M\cap\delta_{G}(Y)|=3$, and $|M^{\prime}\cap\delta_{G}(Y)|=1$. [^2] For a detailed analysis, we refer the reader to the proof of ([6], Theorem 5.4), or ([13], Theorem 13.6). Moving on, we claim that $\delta_{G}(Y)$ is a separating facet-defining cut of $G$. Clearly, $\delta_{G\setminus e}(Y)$ is separating in $G\setminus e$ as it is so in $H$. Thus, as $e\in M^{\prime}$ and $|M^{\prime}\cap\delta_{G}(Y)|=1$, it follows that $\delta_{G}(Y)$ is separating in $G$. To see that $\delta_{G}(Y)$ is facet-defining in $G$, pick perfect matchings $M_{1},\ldots,M_{d}$ whose incidence vectors lie inside the facet $P_{e}(G)$ and are linearly independent; we may pick the first d1d-1 perfect matchings to have intersection $1$ with $\delta_{G}(Y)$, and the last one with intersection $5$, by 3.9. We then swap $M_{d}$ with the perfect matching $M^{\prime}\ni e$ which satisfies $|M^{\prime}\cap\delta(Y)|=1$, in order to obtain $d$ linearly independent vectors $M_{1},\ldots,M_{d-1},M^{\prime}$ in the face $P(G;\delta_{G}(Y))$, implying in turn that $\delta_{G}(Y)$ defines a facet of $P(G)$. Consequently, $(M,\delta_{G}(Y))$ is a desired pair.
@@
In the remaining case, $P(G;C)$ is a (d2)(d-2)-face of $P(G)$, contained in the facet $P_{e}(G)$. Consider now the other facet of $P(G)$ containing $P(G;C)$. It cannot be exposed by a non-negativity inequality, given that $P(G;C)$ is not contained in $\{x:x_{f}=0\}$ for any $f\in E\setminus e$. In particular, the other facet is exposed by a separating facet-defining cut of $G$, say $D$. In particular, $D\setminus e$ defines the same facet as $C\setminus e$, and so $D\setminus e,C\setminus e$ must be equivalent cuts as $G\setminus e$ is a near-brick, by 3.8. Thus, $|M\cap D|=|M\cap C|=3$, implying in turn that $(M,D)$ is a desired pair.
@@
Suppose some (d2)(d-2)-face of $P(G)$ is not exposed by a non-negativity inequality. Take a facet containing this face, which is exposed by a separating facet-defining cut $C=\delta(X)$. Then both $C$-contractions of $G$ are near-bricks, by 3.5, part (3). Given that the polytope $P(G;C)$ has a facet not exposed by a non-negativity inequality, one of the $C$-contractions, say $G/\bar{X}$, must be non-BvN, by 3.3.
We may assume that $G/\bar{X}$ is a non-BvN brick. To argue this, note first that the unique brick of $G/\bar{X}$ is non-BvN, by 3.3. If $G/\bar{X}$ is not this brick, then choose a minimal vertex subset $Y\subset X$ such that $\delta(Y)$ is a tight cut of $G/\bar{X}$. In particular, $P(G;C)\subseteq P(G;\delta(Y))$. As $G$ is a brick, $\delta(Y)$ is not tight in $G$, so $P(G;C)=P(G;\delta(Y))$, therefore $C,\delta(Y)$ are equivalent cuts and $G/Y/\bar{X}$ is bipartite, by 3.8. This, together with our minimal choice of $Y$, implies that $G/\bar{X}/\bar{Y}=G/\bar{Y}$ is a brick. Subsequently, by changing $C=\delta(X)$ and without changing the corresponding facet, if necessary, we may assume $G/\bar{X}$ is a non-BvN brick.
@@
    Suppose $Z\subset X$. Then $D$ is an odd cut of $G/\bar{X}$, so $R,D$ define the same facet of $P(G/\bar{X})$, so $R,D$ are equivalent cuts of $G/\bar{X}$ as $G/\bar{X}$ is a brick, by 3.8. Thus, $|D\cap N|=|R\cap N|=3$, so $|D\cap M|=|D\cap N|=3$, therefore $(M,D)$ is a desired pair.
@@
    Suppose $X\subset Z$. Then $(R,C,D)$ would be a cut triple defining a face triple, thus contradicting 3.10.
@@
    In the remaining case, $X$ and $Z$ cross. Let $I:=\delta(X\cap Z)$ and $U:=\delta(X\cup Z)$. Given that $P(G;C)\cap P(G;D)=P(G;R)\not\subseteq\{x:x_{e}=0\}$ for any $e\in E$, it follows from 3.2 that $x(C)+x(D)=x(I)+x(U)$ and $P(G;R)=P(G;C)\cap P(G;D)=P(G;I)\cap P(G;U)$.
    If $\{P(G;I),P(G;U)\}=\{P(G;R),P(G;C)\}$ or $\{P(G;R),P(G;D)\}$, then $(I,D,U)$ or $(I,C,U)$ would be a cut triple defining a face triple, respectively, thus contradicting 3.10.
    If $\{P(G;I),P(G;U)\}=\{P(G;R),P(G)\}$, then we have $P(G;I)=P(G;R)$ and $P(G;U)=P(G)$. This holds since otherwise $(R,C,U)$ would be a cut triple defining a face triple, thus contradicting 3.10. The equality $P(G;U)=P(G)$ implies that $U$ is a trivial cut as $G$ is a brick, while $P(G;I)=P(G;R)$ implies that $R,I$ are equivalent cuts in the brick $G/\bar{X}$, so $x(I)=x(R)$ for all $x\in P(G)$. Consequently, x(D) = x(I)+x(U)-x(C)=x(R)+1-x(C), so whenever $x(C)=1$, then $x(R)=x(D)$, so $|D\cap M|=|R\cap M|=3$, so $(M,D)$ is a desired pair.
    Otherwise, $\{P(G;I),P(G;U)\}=\{P(G;C),P(G;D)\}$, then we have $P(G;I)=P(G;D)$ and $P(G;U)=P(G;C)$. This holds since otherwise $(R,C,U)$ would be cut triple defining a face triple, thus contradicting 3.10. Now by replacing $D$ with $I$ we fall back to Case 1.
@@ 5 Proof of the Main Theorem
Suppose there is a tight cut $C$. Let $G_{1},G_{2}$ be the $C$-contractions of $G$, both of which are Petersen-free matching-covered graphs. By the induction hypothesis, for each $i\in[2]$, $\operatorname{lin}(P(G_{i}))$ has an integral basis $B^{i}$ that consists solely of the incidence vectors of some perfect matchings of $G_{i}$. Let $B:=B^{1}\odot B^{2}$ which is a subset of $\mathbb{Z}^{E}$ consisting of incidence vectors of some perfect matchings of $G$ that intersect $C$ just once, and is an integral basis for $\operatorname{lin}(P(G;C))$, by 3.1, part (3). Given that $C$ is a tight cut, it follows that $P(G;C)=P$, thus completing the induction step.
@@
Otherwise, $G$ is a non-BvN non-Petersen brick. By the Intersection Theorem, there exist a separating facet-defining cut $C=\delta(X)$ and a perfect matching $M$ such that $|C\cap M|=3$. Let $G_{1}:=G/X$ and $G_{2}:=G/\bar{X}$, both of which are near-bricks, by 3.5, part (3).
Suppose in the first case that both $G_{1},G_{2}$ are Petersen-free. Then by the induction hypothesis, for each $i\in[2]$, $\operatorname{lin}(P(G_{i}))$ has an integral basis $B^{i}$ that consists solely of the incidence vectors of some perfect matchings of $G_{i}$. Let $B^{\prime}:=B^{1}\odot B^{2}$ which is a subset of $\mathbb{Z}^{E}$ consisting of incidence vectors of some perfect matchings of $G$ that intersect $C$ just once, and is an integral basis for $\operatorname{lin}(P(G;C))$, by 3.1, part (3). We claim that $B:=B^{\prime}\cup\{\mathbf{1}_{M}\}$ is an integral basis for $\operatorname{lin}(P)$. Linear independence is clear as $|M\cap C|>1$ while $b(C)=1$ for all $b\in B^{\prime}$. Given that $\dim(P)=\dim(P(G;C))+1$, it therefore follows that $B$ is a basis for $\operatorname{lin}(P)$. It remains to prove that $B$ is an integral basis. To this end, let $z\in\operatorname{lin}(P)\cap\mathbb{Z}^{E}$, and express this vector as a unique linear combination of $B$: $z=\sum_{b\in B}\alpha_{b}b$. We need to show that $\alpha_{b}\in\mathbb{Z}$ for all $b\in B$.
@@
We may assume that the near-brick $G_{2}$ is a Petersen brick. If not, then choose a minimal subset $Z\subset X$ such that $\delta(Z)$ is a tight cut of $G/\bar{X}$. In particular, $P(G;C)\subseteq P(G;\delta(Z))$. As $G$ is a brick, $\delta(Z)$ is not tight in $G$, so $\delta(Z)$ and $C$ define the same facet of $P$, so they are equivalent cuts and $G/Z/\bar{X}$ is bipartite, by 3.8. This, together with our minimal choice of $Z$, implies that $G/\bar{Z}$ is a brick. Subsequently, by changing $C=\delta(X)$ to an equivalent cut if necessary, we may assume $G_{2}$ is a Petersen brick.
@@
To argue that $\delta(Y)$ is facet-defining for $G$, note first that the perfect matching $M$ can be redefined inside $G[X]$, if necessary, such that $|M\cap\delta(Y)|=1$. Secondly, by 3.1, part (4) and 3.9, there exist perfect matchings $M_{1},\ldots,M_{d}$ whose incidence vectors belong to $P(G;C)$ and form a basis for $\operatorname{lin}(P(G;C))$, where $|M_{1}\cap\delta(Y)|=5$, $|M_{i}\cap\delta(Y)|=1$ for $i\in\{2,\ldots,d\}$, and every edge of $G$ belongs to one of $M_{2},\ldots,M_{d}$. By swapping $M_{1}$ with $M$, we obtain $d$ perfect matchings $M,M_{2},\ldots,M_{d}$ whose incidence vectors belong to $P(G;\delta(Y))$ and are linearly independent, because $|M\cap C|=3$ while $|M_{i}\cap C|=1$ for $i\in\{2,\ldots,d\}$. Thus, $P(G;\delta(Y))$ is a facet of $P$.
@@
where each $A_{i}$ is the edge set of some $5$-cycle of $H_{i}$. Let $B:=B_{1}\odot B_{2}$, which clearly consists of some perfect matchings of $G$. By 3.1, part (2), $B$ is a lattice basis for $L$. Furthermore, we claim that the above set equalities imply
@@
$(\subseteq)$ is clear. $(\supseteq)$: Pick $f\in\bar{L}$ such that $f(A_{i})\equiv 0\pmod{2},~\forall i\in[p]$. Then $f=x\odot y$ for $x\in\bar{L}_{1}$ and $y\in\bar{L}_{2}$, where clearly $x(A_{i})\equiv 0\pmod{2},~i=1,\ldots,q$ and $y(A_{i})\equiv 0\pmod{2},~i=q+1,\ldots,p$. Thus, $x\in L_{1},y\in L_{2}$, so $f$ belongs to $L$, by 3.1, part (2). This completes the induction step. ∎
Acknowledgements and declarations
@@
Conflicts of interests/competing interests statement.
The authors have no conflicts of interest nor competing interests to declare that are relevant to the content of this article.